# Testing the cosmological Euler equation: viscosity, equivalence principle, and gravity beyond general relativity

### v2

@@ Abstract
We investigate how the cosmological Euler equation can be tested in the presence of viscous dark matter, violations of the equivalence principle (EP), and modifications of gravity, while relying on minimal theoretical assumptions. Extending the analysis of Castello:2024lhl, we generalize the observable $E_{P}$, which quantifies EP violation, to $\tilde{E}_{P}$, discuss the degeneracy between bulk and shear viscosities and EP-violating effects, and explicitly show that the EP can still be tested in the small-viscosity limit. In addition, we identify a model-independent observable, $C_{\rm vis,0}$, which characterizes the present-day dark matter viscosity and can be measured from relativistic galaxy number counts by cross-correlating two galaxy populations. We perform forecasts for three forthcoming Stage-IV surveys: DESI, Euclid, and SKA Phase 2 (SKA2), and find that $C_{\rm vis,0}$ can be tightly constrained, at the level of $\mathcal{O}(10^{-6})$ or better in all cases. Among these surveys, SKA2 provides the tightest constraint, with a $1\sigma$ uncertainty of $1.08\times 10^{-7}$ on $C_{\rm vis,0}$.
@@ I Introduction
The cosmological Euler equation is a cornerstone of large-scale structure (LSS) formation, as it governs the evolution of velocity perturbations and determines how matter responds to gravitational potentials. Moreover, it encodes the weak equivalence principle, which states that all test particles fall identically in a gravitational field. When testing this relation, it is commonly assumed that dark matter (DM) behaves as a “cold”, pressureless, perfect fluid. This assumption is made both in forecasts (e.g. Euclid:2019clj; Castello:2024lhl) and in the analysis of real data (e.g. Grimm:2025fwc) However, the microscopic nature of dark matter remains unknown, leaving room for more general scenarios when testing gravity on cosmological scales. In this work, we lift the perfect-fluid assumption and allow dark matter to exhibit dissipative properties, focusing on dark matter viscosity.
@@
Deviations from the standard Euler and Poisson equations can lead to scenarios involving EP violation and modified gravity. These modifications can generally be parameterized by several phenomenological free functions that may depend on both time and scale. For instance, deviations in the Euler equation can be captured by an additional friction term $\Theta$ and a fifth-force term $\Gamma$ acting on dark matter (see e.g. Amendola:2003wa; Bonvin:2018ckp). On the other hand, modifications to the Poisson equations are commonly described by an effective gravitational coupling $\mu_{\rm G}$ and a gravitational slip $\eta$ ( see e.g. Pogosian:2010tj; Amendola:2007rr; Amendola:2013qna). The main goal of this paper is to identify the effective combinations of parameters of the Euler equation that can be extracted from LSS observations. As a consequence, we will also find to what extent the parameters that encode the modified gravity and EP violation effects are degenerate with dark matter viscosity. In order to obtain robust conclusions, we remain agnostic about the specific form of the free functions, aiming to cover as broad a class of models as possible.
@@
Viscosity in dark matter may arise from various microscopic processes and can affect cosmic evolution in different ways. At the background level, only the isotropic bulk viscosity contributes, while shear viscosity becomes important at the perturbation level. Previous studies (see e.g. Barbosa:2017ojt; Velten:2012uv; Velten:2013pra; Velten:2014xca) have shown that viscous dark matter can suppress the growth of structure on small scales. We revisit this result and derive the corresponding perturbation equations in a more general setup. Specifically, we focus on how viscosity modifies the perturbed Euler and Einstein equations in the broader context of both EP violation and modified gravity.
@@
Compared to the “standard” way of testing gravity with redshift-space distortions (see e.g. Zheng:2023yco; Sakr:2025kgq), relativistic effects in galaxy number counts Bonvin_2011; Challinor:2011bk provide a powerful probe for testing gravity on large cosmological scales, especially as they offer a way to test gravity without assuming the standard Euler equation Sobral-Blanco:2021cks; Tutusaus:2022cab; Castello:2024jmq. These include corrections such as gravitational redshift, Doppler corrections, (Integrated) Sachs-Wolfe effects, and lensing contributions in the observed galaxy power spectra. Including these corrections becomes increasingly important for upcoming surveys that will be able to detect much larger cosmic volumes, where relativistic effects can no longer be neglected. Forthcoming Stage-IV surveys, such as DESI DESI:2016fyo, Euclid Euclid:2024yrr, and SKA2 Bull:2015lja, will enhance the detectability of relativistic signals and potential deviations from standard gravity. In Castello:2024lhl, a new observable $E_{P}$ was defined as a linear combination of $\Theta$ and $\Gamma$, where $E_{P}\neq 1$ denotes EP violation. It provides a null and model-independent test of EP and is sensitive to scenarios such as coupled quintessence Amendola:1999er In this work, we investigate whether the observable $E_{P}$, or its viscous generalization $\tilde{E}_{P}$, remains a smoking gun for EP violation in the presence of dark matter viscosity.
@@
This work goes beyond previous studies by developing, for the first time, a pipeline to measure the dimensionless viscosity parameter $C_{\rm vis,0}$ at cosmological scales in a significantly more model-independent way. While previous studies have made efforts to constrain dark matter viscosity on both cosmological and smaller astrophysical scales Velten:2012uv; Velten:2013pra; Velten:2014xca; Barbosa:2017ojt; Ashoorioon:2023jwf; AparicioResco:2020shd; Anand:2017wsj; Goswami:2016tsu, they typically rely on strong assumptions, such as the validity of specific parametrizations of structure growth. In contrast, our approach avoids assumptions about the background expansion, galaxy bias functions, or initial conditions of the Universe that determine the shape of the power spectrum. As we will show, despite these minimal assumptions, upcoming Stage-IV surveys will enable precise constraints on viscosity, mainly through the scale-dependent imprint of viscosity on structure growth. This provides a robust test for detecting viscosity, even in the presence of modified gravity and EP violations, without requiring knowledge of their specific physical mechanisms.
@@ II Cosmological Perturbations with Viscous Dark Matter
We consider a perturbed flat Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime and treat the dark matter component as a pressureless fluid with constant bulk and shear viscosities, denoted by $\zeta$ and $\eta_{s}$, respectively. Within the framework of Eckart’s theory of relativistic viscous fluids (cf. Eckart:1940te; Weinberg:1971mx), the energy-momentum tensor of the dark matter fluid is given by:
@@
in which $\theta_{m}=-i\,\mathbf{k}\cdot\mathbf{V}/\mathcal{H}$ is the rescaled dark matter velocity divergence in Fourier space, and $\lambda=\mathcal{H}/k$ characterizes the scale dependence. The perturbed Euler equation agrees with the one presented in Barbosa:2017ojt, for constant viscosities and if $\Phi^{\prime}$ is neglected. In this work, we keep $\Phi^{\prime}$-terms for generality, even if later we find them to be negligible. In addition, from the perturbed Einstein equations, the gravitational potentials are given by
@@
Note that the presence of shear viscosity induces an anisotropic stress ($\Psi\neq\Phi$), which was also noted in Barbosa:2017ojt. This constitutes a degeneracy between shear viscosity and some modified gravity models, further investigated in Barbosa_2018. Degeneracies between modifications of gravity and imperfect dark matter were also studied in AparicioResco:2020shd. In this work, however, we will follow a much more model-independent approach.
@@ III Viscous dark matter, the equivalence principle, and modified gravity
On the other hand, in the scenario of modified gravity and EP violation (see e.g. Amendola:2001rc; Castello:2022uuu; Castello:2024lhl), the perturbed continuity equation remains the same as in Eq. (18), while the Euler equation now becomes
@@ IV Viscous dark matter and galaxy clustering
where $N(\hat{\mathbf{n}},z)$ is the number of galaxies detected in a given pixel, in the direction $\hat{\mathbf{n}}$ and at redshift $z$, and $\bar{N}(z)$ denotes the average number of galaxies per pixel at that redshift. In the linear regime, keeping terms up to order $\lambda$, the observed galaxy number count fluctuation $\Delta$, including relativistic corrections, is given by (see e.g. Bonvin_2011; Challinor_2011; Yoo_2009),
@@
with $s$ denoting the magnification bias arising from the flux-limited nature of surveys and $f^{\rm evol}$ representing the evolution bias, which accounts for changes in the comoving number density due to galaxy formation, evolution, and selection effects. For a more detailed discussion on these biases, see e.g. Maartens:2021dqy.
@@
Once $\Delta$ is obtained, one can construct the observed auto- and cross-power spectra. As in Castello:2024lhl, we consider two galaxy populations, namely a bright (B) and a faint (F) galaxy population. The expressions for $\Delta_{B},\Delta_{F}$ (keeping terms up to the order of $\lambda$) are given by
@@
Here, $P$ denotes the matter power spectrum evaluated at redshift $z=0$, and $B\equiv b_{g,B}b_{g,F}G^{2}(k,z)=f^{2}G^{2}/\beta_{B}\beta_{F}$, where $G(k,z)\equiv\delta_{m}(k,z)/\delta_{m}(k,0)$ is the linear growth factor and $\beta_{T}(k,z)=f(k,z)/b_{g,T}$ with $T=B,F$. The function $S_{g}(k,\mu,z)$ encodes a damping factor $\sigma_{g}$ that accounts for non-linear redshift-space distortions, and here we employ the form suggested in Koda:2013eya; Howlett:2017asw,
@@
The power spectra in Eqs. (50)-(52) are based on the flat-sky approximation. That is, they are derived under the assumption that the separation between the two galaxies in a pair is much smaller than their distance from the observer. A full consideration requires the so-called wide-angle (WA) correction, which is calculated by performing a multipole expansion of the galaxy two-point correlation function in real space and depends on the geometrical configuration. As calculated in Castello:2024lhl, at the order of $\lambda$, the WA term arises only from the cross-power spectra, and it reads
@@
Next, we examine whether $E_{P}$ remains a smoking gun for EP violation in the context of viscous dark matter. Notice that $\tilde{\beta}_{B}$ essentially replaces the standard RSD coefficient $\beta_{B}$, and likewise for $\tilde{\beta}_{F}$. In Castello:2024lhl, the coefficient of the dipole term $i\mu\lambda$ in $P_{\Delta_{B}\Delta_{F}}$ was denoted by $\tau_{1}$, while the coefficient of the octupole term $i\mu^{3}\lambda$ was denoted by $\tau_{2}$. Here, we find that $\tau_{2}=\tilde{\beta}_{F}\tilde{\beta}_{B}\left(\alpha_{F}-\alpha_{B}\right)$ maintains the same structure and, similarly, $\tau_{1}$ reads
@@ V Fisher Analysis
Following the methodology described in Refs. Quartin:2021dmr; Castello:2024lhl, we construct the data covariance matrix in bins of wavenumber $k=|\mathbf{k}|$ and cosine angle $\mu$. The covariance matrix then takes the form of a $2\times 2$ matrix that includes the auto- and cross-power spectra of the bright and faint galaxy populations, together with their shot noise contributions,
@@
Here, $E(z)\equiv H(z)/H_{0}$ is the dimensionless Hubble function, $L_{A}(z)=H_{0}D_{A}(z)$, and $D_{A}(z)$ is the comoving angular diameter distance in the true cosmology. We use the subscript $r$ to refer to the fiducial model. As in Castello:2024lhl, we use the parameter combination $h_{d}\equiv hd$ as a parameter due to the degeneracy between $d$ and $P(k)$ at the linear level. Moreover, we remark that $\mathcal{H}r$ can be expressed using AP parameters, namely $\mathcal{H}r=E_{r}L_{A,r}h_{d}$.
@@
To preserve the independence from the initial conditions of the Universe, we model the shape of the power spectrum using a set of wavebands defined in the first redshift bin. Specifically, we sample $P(k)$ over the range $k\in(0.01,0.12)\,h\,\mathrm{Mpc}^{-1}$ with steps of $\Delta k=0.01\,h\,\mathrm{Mpc}^{-1}$, resulting in a total of twelve waveband parameters. The parameters $d,s_{F},s_{B}$ can be measured independently of galaxy clustering. We can in fact measure $d$ directly with supernovae Ia, and $s_{B,F}$ via number densities as a function of flux (see Appendix H for details). We can use the constraints on these parameters as prior in the Fisher matrix analysis. For the distance parameter $d$, we impose a 3% prior uncertainty, consistent with current constraints from the Hubble diagram and the Pantheon+ supernova sample Brout:2022vxf. The priors on $\alpha_{B}$ and $\alpha_{F}$ are derived by propagating the expected uncertainties on $s_{B}$ and $s_{F}$, as detailed in Sec. VI. Note that in this propagation, we fix $H,r$ to their fiducial values. Moreover, to eliminate the full degeneracy between $P(k)$ and $B$ in the first redshift bin, we fix $B_{z}(z_{1})=1$ and use $B_{z}(z)$ to quantify the ratio of the combination $b_{g,B}\,b_{g,F}\,G_{z}(z)^{2}$ relative to that bin. Finally, to ensure the validity of the perturbative expansion, we adopt the same scale cut procedure as in Castello:2024lhl, namely excluding bins where the $\lambda$-hierarchy breaks down. Specifically, we require that the leading-order terms ($\lambda^{0}$) dominate over the first-order corrections ($\lambda^{1}$), thereby avoiding contamination from higher-order terms. In practice, this condition is enforced by keeping only bins where the Fisher matrix is positive definite, which results in the exclusion of a few low-$k$ bins in a few low-redshift intervals.

### v3

@@ Abstract
We investigate how the cosmological Euler equation can be tested in the presence of viscous dark matter, violations of the equivalence principle (EP), and modifications of gravity, while relying on minimal theoretical assumptions. Extending the analysis of Castello et al. (2025), we generalize the observable $E_{P}$, which quantifies EP violation, to $\tilde{E}_{P}$, discuss the degeneracy between bulk and shear viscosities and EP-violating effects, and explicitly show that the EP can still be tested in the small-viscosity limit. In addition, we identify a model-independent observable, $C_{\rm vis,0}$, which characterizes the present-day dark matter viscosity and can be measured from relativistic galaxy number counts by cross-correlating two galaxy populations. We perform forecasts for three forthcoming Stage-IV surveys: DESI, Euclid, and SKA Phase 2 (SKA2), and find that $C_{\rm vis,0}$ can be tightly constrained, at the level of $\mathcal{O}(10^{-6})$ or better in all cases. Among these surveys, SKA2 provides the tightest constraint, with a $1\sigma$ uncertainty of $1.08\times 10^{-7}$ on $C_{\rm vis,0}$.
@@ I Introduction
The cosmological Euler equation is a cornerstone of large-scale structure (LSS) formation, as it governs the evolution of velocity perturbations and determines how matter responds to gravitational potentials. Moreover, it encodes the weak equivalence principle, which states that all test particles fall identically in a gravitational field. When testing this relation, it is commonly assumed that dark matter (DM) behaves as a “cold”, pressureless, perfect fluid. This assumption is made both in forecasts (e.g. Blanchard and others (2020); Castello et al. (2025)) and in the analysis of real data (e.g. Grimm et al. (2025)) However, the microscopic nature of dark matter remains unknown, leaving room for more general scenarios when testing gravity on cosmological scales. In this work, we lift the perfect-fluid assumption and allow dark matter to exhibit dissipative properties, focusing on dark matter viscosity.
@@
Deviations from the standard Euler and Poisson equations can lead to scenarios involving EP violation and modified gravity. These modifications can generally be parameterized by several phenomenological free functions that may depend on both time and scale. For instance, deviations in the Euler equation can be captured by an additional friction term $\Theta$ and a fifth-force term $\Gamma$ acting on dark matter (see e.g. Amendola (2004); Bonvin and Fleury (2018)). On the other hand, modifications to the Poisson equations are commonly described by an effective gravitational coupling $\mu_{\rm G}$ and a gravitational slip $\eta$ ( see e.g. Pogosian et al. (2010); Amendola et al. (2008, 2014)). The main goal of this paper is to identify the effective combinations of parameters of the Euler equation that can be extracted from LSS observations. As a consequence, we will also find to what extent the parameters that encode the modified gravity and EP violation effects are degenerate with dark matter viscosity. In order to obtain robust conclusions, we remain agnostic about the specific form of the free functions, aiming to cover as broad a class of models as possible.
@@
Viscosity in dark matter may arise from various microscopic processes and can affect cosmic evolution in different ways. At the background level, only the isotropic bulk viscosity contributes, while shear viscosity becomes important at the perturbation level. Previous studies (see e.g. Barbosa et al. (2017); Velten and Schwarz (2012); Velten et al. (2013, 2014)) have shown that viscous dark matter can suppress the growth of structure on small scales. We revisit this result and derive the corresponding perturbation equations in a more general setup. Specifically, we focus on how viscosity modifies the perturbed Euler and Einstein equations in the broader context of both EP violation and modified gravity.
@@
Compared to the “standard” way of testing gravity with redshift-space distortions (see e.g. Zheng et al. (2024); Sakr et al. (2025)), relativistic effects in galaxy number counts Bonvin and Durrer (2011); Challinor and Lewis (2011b) provide a powerful probe for testing gravity on large cosmological scales, especially as they offer a way to test gravity without assuming the standard Euler equation Sobral-Blanco and Bonvin (2021); Tutusaus et al. (2023); Castello et al. (2024b). These include corrections such as gravitational redshift, Doppler corrections, (Integrated) Sachs-Wolfe effects, and lensing contributions in the observed galaxy power spectra. Including these corrections becomes increasingly important for upcoming surveys that will be able to detect much larger cosmic volumes, where relativistic effects can no longer be neglected. Forthcoming Stage-IV surveys, such as DESI Aghamousa and others (2016), Euclid Mellier and others (2025), and SKA2 Bull (2016), will enhance the detectability of relativistic signals and potential deviations from standard gravity. In Castello et al. (2025), a new observable $E_{P}$ was defined as a linear combination of $\Theta$ and $\Gamma$, where $E_{P}\neq 1$ denotes EP violation. It provides a null and model-independent test of EP and is sensitive to scenarios such as coupled quintessence Amendola (2000). In this work, we investigate whether the observable $E_{P}$, or its viscous generalization $\tilde{E}_{P}$, remains a smoking gun for EP violation in the presence of dark matter viscosity.
@@
This work goes beyond previous studies by developing, for the first time, a pipeline to measure the dimensionless viscosity parameter $C_{\rm vis,0}$ at cosmological scales in a significantly more model-independent way. While previous studies have made efforts to constrain dark matter viscosity on both cosmological and smaller astrophysical scales Velten and Schwarz (2012); Velten et al. (2013, 2014); Barbosa et al. (2017); Ashoorioon and Davari (2023); Aparicio Resco and Maroto (2021); Anand et al. (2017); Goswami et al. (2017), they typically rely on strong assumptions, such as the validity of specific parametrizations of structure growth. In contrast, our approach avoids assumptions about the background expansion, galaxy bias functions, or initial conditions of the Universe that determine the shape of the power spectrum. As we will show, despite these minimal assumptions, upcoming Stage-IV surveys will enable precise constraints on viscosity, mainly through the scale-dependent imprint of viscosity on structure growth. This provides a robust test for detecting viscosity, even in the presence of modified gravity and EP violations, without requiring knowledge of their specific physical mechanisms.
@@ II Cosmological Perturbations with Viscous Dark Matter
We consider a perturbed flat Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime and treat the dark matter component as a pressureless fluid with constant bulk and shear viscosities, denoted by $\zeta$ and $\eta_{s}$, respectively. Within the framework of Eckart’s theory of relativistic viscous fluids (cf. Eckart (1940); Weinberg (1971)), the energy-momentum tensor of the dark matter fluid is given by:
@@
in which $\theta_{m}=-i\,\mathbf{k}\cdot\mathbf{V}/\mathcal{H}$ is the rescaled dark matter velocity divergence in Fourier space, and $\lambda=\mathcal{H}/k$ characterizes the scale dependence. The perturbed Euler equation agrees with the one presented in Barbosa et al. (2017), for constant viscosities and if $\Phi^{\prime}$ is neglected. In this work, we keep $\Phi^{\prime}$-terms for generality, even if later we find them to be negligible. In addition, from the perturbed Einstein equations, the gravitational potentials are given by
@@
Note that the presence of shear viscosity induces an anisotropic stress ($\Psi\neq\Phi$), which was also noted in Barbosa et al. (2017). This constitutes a degeneracy between shear viscosity and some modified gravity models, further investigated in Barbosa et al. (2018). Degeneracies between modifications of gravity and imperfect dark matter were also studied in Aparicio Resco and Maroto (2021). In this work, however, we will follow a much more model-independent approach.
@@ III Viscous dark matter, the equivalence principle, and modified gravity
On the other hand, in the scenario of modified gravity and EP violation (see e.g. Amendola and Tocchini-Valentini (2002); Castello et al. (2022, 2025)), the perturbed continuity equation remains the same as in Eq. (18), while the Euler equation now becomes
@@ IV Viscous dark matter and galaxy clustering
where $N(\hat{\mathbf{n}},z)$ is the number of galaxies detected in a given pixel, in the direction $\hat{\mathbf{n}}$ and at redshift $z$, and $\bar{N}(z)$ denotes the average number of galaxies per pixel at that redshift. In the linear regime, keeping terms up to order $\lambda$, the observed galaxy number count fluctuation $\Delta$, including relativistic corrections, is given by (see e.g. Bonvin and Durrer (2011); Challinor and Lewis (2011a); Yoo et al. (2009)),
@@
with $s$ denoting the magnification bias arising from the flux-limited nature of surveys and $f^{\rm evol}$ representing the evolution bias, which accounts for changes in the comoving number density due to galaxy formation, evolution, and selection effects. For a more detailed discussion on these biases, see e.g. Maartens et al. (2021).
@@
Once $\Delta$ is obtained, one can construct the observed auto- and cross-power spectra. As in Castello et al. (2025), we consider two galaxy populations, namely a bright (B) and a faint (F) galaxy population. The expressions for $\Delta_{B},\Delta_{F}$ (keeping terms up to the order of $\lambda$) are given by
@@
Here, $P$ denotes the matter power spectrum evaluated at redshift $z=0$, and $B\equiv b_{g,B}b_{g,F}G^{2}(k,z)=f^{2}G^{2}/\beta_{B}\beta_{F}$, where $G(k,z)\equiv\delta_{m}(k,z)/\delta_{m}(k,0)$ is the linear growth factor and $\beta_{T}(k,z)=f(k,z)/b_{g,T}$ with $T=B,F$. The function $S_{g}(k,\mu,z)$ encodes a damping factor $\sigma_{g}$ that accounts for non-linear redshift-space distortions, and here we employ the form suggested in Koda et al. (2014); Howlett et al. (2017),
@@
The power spectra in Eqs. (50)-(52) are based on the flat-sky approximation. That is, they are derived under the assumption that the separation between the two galaxies in a pair is much smaller than their distance from the observer. A full consideration requires the so-called wide-angle (WA) correction, which is calculated by performing a multipole expansion of the galaxy two-point correlation function in real space and depends on the geometrical configuration. As calculated in Castello et al. (2025), at the order of $\lambda$, the WA term arises only from the cross-power spectra, and it reads
@@
Next, we examine whether $E_{P}$ remains a smoking gun for EP violation in the context of viscous dark matter. Notice that $\tilde{\beta}_{B}$ essentially replaces the standard RSD coefficient $\beta_{B}$, and likewise for $\tilde{\beta}_{F}$. In Castello et al. (2025), the coefficient of the dipole term $i\mu\lambda$ in $P_{\Delta_{B}\Delta_{F}}$ was denoted by $\tau_{1}$, while the coefficient of the octupole term $i\mu^{3}\lambda$ was denoted by $\tau_{2}$. Here, we find that $\tau_{2}=\tilde{\beta}_{F}\tilde{\beta}_{B}\left(\alpha_{F}-\alpha_{B}\right)$ maintains the same structure and, similarly, $\tau_{1}$ reads
@@ V Fisher Analysis
Following the methodology described in Refs. Quartin et al. (2022); Castello et al. (2025), we construct the data covariance matrix in bins of wavenumber $k=|\mathbf{k}|$ and cosine angle $\mu$. The covariance matrix then takes the form of a $2\times 2$ matrix that includes the auto- and cross-power spectra of the bright and faint galaxy populations, together with their shot noise contributions,
@@
In principle, our $\mu$-binning or a sufficiently high multipole expansion (often adopted in previous analyses) contain the same information. However, the $\mu$-binning leaves us more flexibility in experimenting with different number of bins without introducing very high multipoles that likely will not be measured effectively in future surveys. Secondly, the multipole expansion is usually adopted to justify discarding the higher multipoles on the ground that they contain less and less information. This is of course true in most cases, but in the spirit of our model-independent approach we chose not to enforce this additional assumption.
@@
Here, $E(z)\equiv H(z)/H_{0}$ is the dimensionless Hubble function, $L_{A}(z)=H_{0}D_{A}(z)$, and $D_{A}(z)$ is the comoving angular diameter distance in the true cosmology. We use the subscript $r$ to refer to the fiducial model. As in Castello et al. (2025), we use the parameter combination $h_{d}\equiv hd$ as a parameter due to the degeneracy between $d$ and $P(k)$ at the linear level. Moreover, we remark that $\mathcal{H}r$ can be expressed using AP parameters, namely $\mathcal{H}r=E_{r}L_{A,r}h_{d}$.
@@
For convenience, we summarize in Table 2 the primary definitions of these parameters as introduced in the text. To preserve independence from the initial conditions of the Universe, we model the shape of the power spectrum using a set of wavebands defined in the first redshift bin. Specifically, we sample $P(k)$ over the range $k\in(0.01,0.12)\,h\,\mathrm{Mpc}^{-1}$ with steps of $\Delta k=0.01\,h\,\mathrm{Mpc}^{-1}$, resulting in a total of twelve waveband parameters. The parameters $d,s_{F},s_{B}$ can be measured independently of galaxy clustering. We can in fact measure $d$ directly with supernovae Ia, and $s_{B,F}$ via number densities as a function of flux (see Appendix H for details). We can use the constraints on these parameters as prior in the Fisher matrix analysis. For the distance parameter $d$, we impose a 3% prior uncertainty, consistent with current constraints from the Hubble diagram and the Pantheon+ supernova sample Brout and others (2022). The priors on $\alpha_{B}$ and $\alpha_{F}$ are derived by propagating the expected uncertainties on $s_{B}$ and $s_{F}$, as detailed in Sec. VI. Note that in this propagation, we fix $H,r$ to their fiducial values. Moreover, to eliminate the full degeneracy between $P(k)$ and $B$ in the first redshift bin, we fix $B_{z}(z_{1})=1$ and use $B_{z}(z)$ to quantify the ratio of the combination $b_{g,B}\,b_{g,F}\,G_{z}(z)^{2}$ relative to that bin. Finally, to ensure the validity of the perturbative expansion, we adopt the same scale cut procedure as in Castello et al. (2025), namely excluding bins where the $\lambda$-hierarchy breaks down. Specifically, we require that the leading-order terms ($\lambda^{0}$) dominate over the first-order corrections ($\lambda^{1}$), thereby avoiding contamination from higher-order terms. In practice, this condition is enforced by keeping only bins where the Fisher matrix is positive definite, which results in the exclusion of a few low-$k$ bins in a few low-redshift intervals.
parameters description definitions in the text
$\tilde{\beta}_{\mathrm{B},z},\tilde{\beta}_{\mathrm{F},z}$ purely time-dependent component of $\tilde{\beta}_{B}$ and $\tilde{\beta}_{F}$ Eqs. (43) and (44)
$\alpha_{B},\alpha_{F}$ encoding the magnification and evolution bias Eq. (30)
$h_{d},d$ AP effect parameters Eq. (63) and the paragraph below
$B_{z}$ purely time-dependent component of $B$ Eq. (46)
$\sigma_{g}$ non-linear RSD damping factor Eq. (42)
$\tilde{E}_{P,z}$ parameter indicating EP violation Eqs. (35) and (45)
$n_{n}$ shot noise amplitude Eq. (58)
Table 2: List of free parameters varied in each redshift bin, along with their definitions in the text.